OpenStudy (anonymous):

Jake's band played songs at three different parties. At the first party they played four-fifths as long as they did at the second party, which was one-half as long as the third party. If they played for 6 minutes at the first party, how many minutes did they play at the third party? 2 minutes four and four-fifths minutes eight and three-fifths minutes fifteen minutes

OpenStudy (anonymous):

@vipergirl81 ...

OpenStudy (anonymous):

So to make it all a bit simpler, you represent the first party as 'a', second party as 'b' and third party as 'c'. a=4/5b c = 1/2 b since a=6, 6=4/5(b) b=7.5 c=2(7.5) c=15 So they played 15 minutes at the third party.

OpenStudy (anonymous):

Sorry it took so long. I got messed up in the middle and had to retype everything.

OpenStudy (anonymous):

@vipergirl81 thank you so much and thanks for the explaintion too!

OpenStudy (anonymous):

no problem!